OpenStudy (anonymous):

Write an equation of the line that passes through (4, 6) and is parallel to the line .

OpenStudy (anonymous):

parallel to what??

OpenStudy (anonymous):

divu.mkr has a point what line are we making a parallel line to?

OpenStudy (anonymous):

y=3x+4

OpenStudy (anonymous):

ok first put in point-slope form y-6=3(x-4) distribute y-6=3x-12 add six y=3x-6

OpenStudy (anonymous):

okay i got that answer i just had to make sure it was right!

OpenStudy (anonymous):

ok well your welcome